OpenStudy (anonymous):

Solve for x: 3x − 24 = 81

OpenStudy (anonymous):

Ok.. what is your understanding of this Q so far?

OpenStudy (anonymous):

Recall you need to isolate your x-variable.

OpenStudy (anonymous):

First add 24 to both sides. 3x-24(+24)=81(+24) 3x=105 Now divide both sides by three 3x/3= x 105/3=35 So your answer is x=35

OpenStudy (anonymous):

Please don't give out direct answers :) @alexistheking777 Instead, try engaging the asker so they can learn better and understand where your answer is coming from. As the saying goes… \(\\\color{darkblue}{Give~a~man~a~fish,~he~eats~for~a~day.}\) \(\\\color{pink}{Teach~a~man~to~fish,~he~eats~for~a~lifetime.}\)

OpenStudy (anonymous):

You didn't entirely give the direct answer, but it would've been better to engage w/ the asker :)